# jina-reranker-v3: Last but Not Late Interaction for Document Reranking

### v2

@@ Abstract
jina-reranker-v3 is a 0.6B parameter multilingual document reranker that introduces a novel last but not late interaction. Unlike late interaction models such as ColBERT that perform separate encoding followed by multi-vector matching, our approach conducts causal self-attention between query and documents within the same context window, enabling rich cross-document interactions before extracting contextual embeddings from the last token of each document. This compact architecture achieves state-of-the-art BEIR performance with 61.94 nDCG@10 while being $10\times$ smaller than generative listwise rerankers.
@@ 1 Introduction
Neural document reranking faces a fundamental efficiency-effectiveness tradeoff. Cross-encoders achieve strong performance through joint query-document processing but require separate forward passes for each pair, while embedding models enable efficient similarity computation but lose fine-grained interaction signals. Recent innovations have attempted to bridge this gap through different interaction approaches. Late interaction models like ColBERT (Khattab and Zaharia, 2020) and their variants (Liu et al., 2024; Chaffin and Sourty, 2025) separately encode queries and documents into multi-vector representations, then perform interaction through token-level similarity operations.
We introduce jina-reranker-v3, which proposes a novel last but not late interaction (LBNL) that fundamentally differs from existing approaches. While late interaction models delay attention interactions until after separate encoding, our method performs causal query-document self-attention within the transformer context window first, enabling documents to attend to each other and establish cross-document relationships before extracting contextual embeddings from the last token of each document. This cross-document interaction capability is unavailable in separate encoding approaches and represents the core innovation of our method. Unlike late interaction models that perform interactions after encoding, we enable interactions during encoding, making our approach not late.
Our evaluation shows significant improvements across diverse benchmarks. jina-reranker-v3 achieves 61.94 nDCG@10 on BEIR, representing the highest score among all evaluated rerankers and a 4.88% improvement over our previous jina-reranker-v2. The model excels particularly in multi-hop retrieval with HotpotQA reaching 78.56, fact verification achieving 93.95 on FEVER, competitive multilingual performance across 18 languages at 66.50 on MIRACL, and specialized domain coverage reaching 63.28 on code retrieval.
@@ 2 Related Work
Late interaction models represent a significant approach that balances efficiency with expressiveness. ColBERT (Khattab and Zaharia, 2020) exemplifies this approach by independently encoding queries and documents into multi-vector representations, then computing similarity through MaxSim operations over token-level embeddings. This design enables pre-computation of document representations while preserving fine-grained matching signals. Recent developments have expanded this approach: analysis of matching mechanisms and token pruning strategies (Liu et al., 2024) provides theoretical foundations, LITE (Ji et al., 2024) introduces learnable late interactions, and Jina-ColBERT-v2 (Zhang et al., 2024) extends the approach to multilingual settings. PyLate (Chaffin and Sourty, 2025) provides flexible frameworks for training and deployment of such models. The late chunking method (Günther et al., 2024) processes complete documents through transformers before applying chunking boundaries, extracting chunk-level embeddings that preserve contextual relationships. This approach demonstrates how leveraging broader document context can improve embedding quality, though it focuses primarily on retrieval rather than reranking applications.
@@ 3.1 Architecture
[…]
Our architecture addresses fundamental limitations in existing interactions. ColBERT (Khattab and Zaharia, 2020) achieves efficiency through separate encoding followed by multi-vector interaction, but cannot capture early query-document interactions during encoding or enable cross-document interactions within the attention mechanism. Our approach enables causal self-attention interaction within the transformer architecture, allowing documents to attend to each other and the query simultaneously, then extracts contextually-informed embeddings from designated token positions.
The core innovation lies in enabling cross-document interactions during encoding: instead of delaying interaction until after separate encoding as in late interaction models, we process all documents and the query simultaneously within shared context windows. This allows each document to attend to other documents and observe their content, enabling contextual embeddings that capture not just query-document relevance but also inter-document relationships and comparative context. Such cross-document interactions are impossible in separate encoding approaches and represent a fundamental advancement in reranking architecture.
@@ 3.2 Prompt Template
<|im_start|>assistant
<think></think>
Table 1: Complete prompt template structure used by jina-reranker-v3. Special tokens <|doc_emb|> and <|query_emb|> mark positions for embedding extraction from transformer hidden states. The empty <think></think> tag serves as a placeholder without reasoning content.
The template strategically places the query both at the beginning for instructions and at the end for final attention, sandwiching all documents in between. This design enables the final query position to attend to all preceding documents through causal attention while maintaining clear task instructions. Special tokens <|doc_emb|> after each document and <|query_emb|> after the final query mark specific positions for embedding extraction from transformer hidden states. While Qwen3 supports thinking capabilities, our implementation uses an empty <think></think> placeholder since the model is not trained with reasoning traces for document reranking tasks.
@@ 4 Training
$\ell_{\mathrm{disperse}}=\frac{1}{N}\log\frac{1}{K+1}\left(e^{s(\mathbf{d}_{i}^{+},\mathbf{d}_{i,k}^{-})/\tau}+\sum_{k=1}^{K-1}e^{s(\mathbf{d}_{i,k}^{-},\mathbf{d}_{i,k+1}^{-})/\tau}\right)$ (3)
@@
Stage 2: Context and Hard Negative Mining. This stage combines context scaling and comprehensive robustness optimization. Context scaling extends sequence lengths to 8,192 tokens through configurations like mldr-8192 for long documents using multilingual MLDR datasets, jina-crosslingual for enhanced multilingual capabilities, and biomed for medical domains. Simultaneously, cross-system hard negative mining ensures robustness through specialized optimizations including jina-en-v2 for English performance, miracl-v2 for multilingual retrieval, cornstack-v2 for code understanding, and context-chunk-v3 for long-document processing. Training systematically mines hard negatives across multiple retrieval systems including BGE, Jina, GTE, and E5-Large with up to 25 negatives per query and very low temperature of 0.05, using key datasets including MS-MARCO Campos et al. (2016), mMARCO Bonifacio et al. (2021), and domain-specific synthetic question-answer pairs.
@@ 5 Evaluation
We evaluate jina-reranker-v3 across diverse multilingual retrieval benchmarks to demonstrate its architectural innovations and training efficiency. Our comprehensive assessment reveals how the dual embedding architecture with LBNL interaction enables superior performance density, achieving state-of-the-art results with remarkable parameter efficiency.
@@ 5.1 Experimental Setup
Our evaluation spans four challenging benchmarks that test different aspects of ranking capability. BEIR Thakur et al. (2021) represents the gold standard for English retrieval evaluation, encompassing 13 heterogeneous tasks from question answering on Natural Questions to fact verification on FEVER, testing the model’s ability to generalize across domains without task-specific optimization. MIRACL Zhang et al. (2023) pushes multilingual boundaries with 18 languages spanning diverse linguistic families, from Arabic and Chinese to Finnish and Thai, requiring deep cross-lingual understanding. MKQA Longpre et al. (2021) specifically challenges cross-lingual question answering capabilities, while CoIR Li et al. (2025) focuses on the specialized domain of code retrieval, where semantic understanding of programming constructs becomes crucial.
We establish rigorous baselines spanning both retrieval approaches. First-stage dense retrievers include jina-embeddings-v3, providing the foundation top-100 candidates that all rerankers process. Second-stage rerankers encompass our previous jina-reranker-v2, the multilingual bge-reranker-v2-m3, the mxbai-rerank variants at different scales, and Qwen3-Reranker-0.6B and Qwen3-Reranker-4B models. This comprehensive comparison enables analysis across both architectural approaches and parameter scaling effects.
All evaluations follow identical protocols using nDCG@10 as the primary metric, with our own controlled runs ensuring fair comparison. The consistent use of jina-embeddings-v3 retrieval candidates eliminates retrieval variance and isolates reranking effectiveness.
@@ 5.2 Overall Performance Across Benchmarks
Table 2 demonstrates jina-reranker-v3’s exceptional performance density across diverse evaluation scenarios. On BEIR, our model achieves the highest score among all rerankers at 61.94, establishing new state-of-the-art performance for English retrieval. This represents a 4.88% improvement over our previous jina-reranker-v2 at 57.06, directly attributable to LBNL interaction mechanism where query and document embeddings are extracted from shared forward passes rather than separate encoding pipelines.
Parameter efficiency analysis reveals striking advantages compared to larger alternatives. Against the 1.5B parameter mxbai-rerank-large-v2, jina-reranker-v3 achieves superior BEIR performance with 61.94 versus 61.44 using $2.5\times$ fewer parameters, while providing specialized domain coverage unavailable in competing models reaching 63.28 on CoIR. This efficiency derives from architectural innovations: Qwen3’s optimized transformer backbone combined with our specialized 256-dimensional projector network that concentrates ranking signals without requiring massive parameter scaling.
Multilingual evaluation reveals strong cross-lingual capabilities despite the model’s compact architecture. The 66.50 score on MIRACL, while 2.82 points below the multilingual-specialized bge-reranker-v2-m3 at 69.32, demonstrates effective knowledge transfer from our progressive training methodology. The 67.84 MKQA performance closely approaches jina-reranker-v2’s 67.90, indicating that architectural sophistication can partially offset parameter differences in multilingual scenarios.
@@
jina-embeddings-v2-base-code - - - - 56.90
@@
jina-reranker-v3 0.6B 61.94 66.50 67.84 63.28
jina-reranker-v2 0.3B 57.06 63.65 67.90 56.14
jina-reranker-m0 2.4B 58.95 66.75 68.19 63.55
bge-reranker-v2-m3 0.6B 56.51 69.32 67.88 35.97
mxbai-rerank-base-v2 0.5B 58.40 55.32 64.24 -
mxbai-rerank-large-v2 1.5B 61.44 57.94 67.06 -
Qwen3-Reranker-0.6B 0.6B 56.28 57.70 65.34 -
Qwen3-Reranker-4B 4.0B 61.16 67.52 - -
Table 2: Evaluation results (nDCG@10 in percent) for reranking models. All scores are our runs based on the retrieval top-100 results from the first row.
@@ 5.3 English Retrieval Performance on BEIR
Table 3 provides granular analysis across BEIR’s heterogeneous tasks, revealing specific architectural advantages. The model achieves consistent excellence across diverse reasoning tasks, with particularly strong performance on complex multi-hop reasoning reaching 78.56 on HotpotQA and fact verification achieving 93.95 on FEVER. These results highlight how LBNL interaction enables sophisticated query-document self-attention during encoding, capturing evidence relationships that separate encoding approaches miss.
Within the same scale category, jina-reranker-v3 reveals significant advantages. Against bge-reranker-v2-m3 with the same 0.6B parameters, jina-reranker-v3 delivers a substantial 5.43% improvement from 56.51 to 61.94, demonstrating architectural innovation over simple parameter scaling. The specialized 256-dimensional projector network effectively concentrates ranking signals while preserving contextual representations from the Qwen3 backbone. Remarkably, our model surpasses mxbai-rerank-large-v2’s 61.44 performance while using $2.5\times$ fewer parameters, establishing that sophisticated architecture can surpass brute-force scaling approaches.
Since jina-reranker-v3 processes all documents simultaneously in a listwise manner within shared context windows, we investigate the sensitivity to document ordering. We evaluate three variants: documents ordered by descending relevance scores (D), ascending scores (A), and random permutation (R). The results show modest variations across orderings, with random ordering (R) achieving the highest average of 62.54, followed by descending (D) at 61.94 and ascending (A) at 61.52. While the differences are not conclusive, this analysis reveals that the LBNL interaction maintains relatively stable performance across different input orderings, suggesting robust self-attention mechanisms that can effectively process documents regardless of their initial arrangement.
The model’s dominance extends particularly to question-answering scenarios, where Natural Questions achieves 74.30 and argumentative retrieval on ArguAna reaches 77.39, showcasing the benefit of contextual embeddings. These tasks require understanding complex query intent and matching it against nuanced document semantics, precisely the scenario where our dual embedding extraction approach provides maximum advantage over traditional cross-encoder scoring.
@@
jina-reranker-v3 (D) 0.6B 61.94 84.32 37.59 74.30 78.56 48.90 77.39 32.11 47.70 23.08 93.95 40.77 75.98 90.60
jina-reranker-v3 (A) 0.6B 61.52 85.78 39.53 72.50 77.39 51.15 69.09 31.46 48.14 23.90 93.40 40.83 76.53 90.10
jina-reranker-v3 (R) 0.6B 62.54 86.39 39.66 73.14 77.91 51.79 76.02 31.29 48.22 24.27 93.69 42.83 77.55 90.31
@@ 5.4 Multilingual Performance on MIRACL
Table 4: Multi-lingual retrieval performance on the MIRACL dev set (measured by nDCG@10).
@@ 6 Conclusion
We present jina-reranker-v3, a 0.6B parameter multilingual document reranker that introduces last but not late interaction for efficient document reranking. Our approach enables cross-document interactions during encoding by processing queries and multiple documents simultaneously within shared context windows, then extracting contextual embeddings from designated special token positions.
By effectively adapting modern long-context language models for document reranking, jina-reranker-v3 demonstrates that discriminative methods can bridge the efficiency-effectiveness gap while maintaining significant computational advantages over generative alternatives. The results show strong performance across diverse benchmarks, with particular excellence in complex retrieval tasks and multilingual scenarios.

### v4

# jina-reranker-v3: Last but Not Late Interaction for Listwise Document Reranking

@@ Abstract
jina-reranker-v3 is a 0.6B-parameter multilingual listwise reranker that introduces a novel last but not late interaction. Unlike late interaction models like ColBERT that encode documents separately before multi-vector matching, our approach applies causal attention between the query and all candidate documents in the same context window, enabling rich interactions before extracting contextual embeddings from each document’s final token. The new model achieves state-of-the-art BEIR performance with 61.85 nDCG@10 while being significantly smaller than other models with comparable performance.
@@ 1 Introduction
Neural document retrieval faces a fundamental efficiency-effectiveness tradeoff. Cross-encoders achieve strong performance through joint query-document processing but require separate forward passes for each pair, while embedding models enable efficient similarity computation but lose fine-grained interaction signals. Recent models have attempted to bridge this gap through different interaction approaches. Late interaction models like ColBERT (Khattab and Zaharia, 2020) and their variants (Liu et al., 2024; Jha et al., 2024) separately encode queries and documents into multi-vector representations, then perform interaction through token-level similarity operations.
We introduce jina-reranker-v3, which features a novel last but not late interaction (LBNL) that takes a fundamentally different approach from existing methods. While late interaction models delay attention until after encoding documents separately, our method applies causal attention between the query and all documents within the context window, enabling cross-document interactions before extracting contextual embeddings from each document’s last token. Unlike late interaction models that interact after encoding, we enable interactions during encoding—making our approach not late. This “listwise” processing is not possible with separate encoding or bi-encoder approaches and represents our core innovation.
Evaluation shows jina-reranker-v3 achieves 61.85 nDCG@10 on BEIR (Thakur et al., 2021), representing the highest score among all evaluated rerankers and a 4.79% improvement over our previous jina-reranker-v2. The model excels particularly in multi-hop retrieval with HotpotQA reaching 78.58, fact verification achieving 94.01 on FEVER, competitive multilingual performance across 18 languages at 66.83 on MIRACL (Zhang et al., 2023a) and crosslingual retrieval with 67.92 Recall@10 on MKQA (Longpre et al., 2020) across 26 languages, and code retrieval reaching 70.64 on CoIR (Li et al., 2024).
@@ 2 Related Work
Late interaction models represent a significant approach that balances efficiency with expressiveness. ColBERT (Khattab and Zaharia, 2020) exemplifies this approach by independently encoding queries and documents into multi-vector representations, then computing similarity through MaxSim operations over token-level embeddings. This design enables pre-computation of document representations while preserving fine-grained matching signals. Recent developments have expanded this approach: analysis of matching mechanisms and token pruning strategies (Liu et al., 2024) provides theoretical foundations, LITE (Ji et al., 2024) introduces learnable late interactions, and Jina-ColBERT-v2 (Jha et al., 2024) extends the approach to multilingual settings. PyLate (Chaffin and Sourty, 2025) provides flexible frameworks for training and deployment of such models. The late chunking method (Günther et al., 2024) processes complete documents through transformers before applying chunking boundaries, extracting chunk-level embeddings that preserve contextual relationships. This approach demonstrates how leveraging broader document context can improve embedding quality, though it focuses primarily on retrieval rather than reranking applications.
@@ 3.1 Architecture
Figure 1 illustrates our architecture that addresses fundamental limitations in existing interactions. ColBERT (Khattab and Zaharia, 2020) achieves efficiency through separate encoding followed by multi-vector interaction, but cannot capture early query-document interactions during encoding or enable cross-document interactions within the attention mechanism.
Our LBNL approach enables causal self-attention interaction within the transformer architecture: instead of delaying interaction until after separate encoding as in late interaction models, we process all documents and the query simultaneously within shared context windows. This allows each document to attend to other documents and observe their content, enabling contextual embeddings that capture not just query-document relevance but also inter-document relationships and comparative context. Such cross-document interactions are impossible in separate encoding approaches and represent a fundamental advancement in jina-reranker-v3 architecture.
@@ 3.2 Prompt Template
Table 1: Complete prompt template structure used by jina-reranker-v3. Special tokens <|doc_emb|> and <|query_emb|> mark positions for embedding extraction from transformer hidden states.
The template strategically places the query both at the beginning for instructions and at the end for final attention, sandwiching all documents in between. This design enables the final query position to attend to all preceding documents through causal attention while maintaining clear task instructions. Special tokens <|doc_emb|> after each document and <|query_emb|> after the final query mark specific positions for embedding extraction from transformer hidden states.
@@ 4.1 Loss Functions
$\ell_{\mathrm{disperse}}=\frac{1}{N}\sum_{i=1}^{N}\log\frac{1}{K}\sum_{k=1}^{K}\left(e^{s(\mathbf{d}_{i}^{+},\mathbf{d}_{i,k}^{-})/\tau}+\sum_{k^{\prime}=k}^{K-1}e^{s(\mathbf{d}_{i,k}^{-},\mathbf{d}_{i,k^{\prime}+1}^{-})/\tau}\right)$ (3)
@@ 4.2 Multi-Stage Training
Stage 2: Context and Hard Negative Mining. This stage combines context extension and comprehensive robustness optimization. Context extension is implemented in two ways: (1) extending individual document length to 8,192 tokens through datasets like MLDR (Chen et al., 2024) for long-document understanding, and (2) increasing the number of negative documents from 15 to 45 per query while maintaining the total sequence length under 131K tokens. Simultaneously, cross-system hard negative mining ensures robustness through specialized optimizations including jina-en-v2 for English performance, miracl-v2 for multilingual retrieval, cornstack-v2 for code understanding, and context-chunk-v3 for long-document processing. Training systematically mines hard negatives across multiple retrieval systems including BGE, Jina, GTE, and E5-Large with up to 25 negatives per query and very low temperature of 0.05, using key datasets including MS-MARCO Campos et al. (2016), mMARCO Bonifacio et al. (2021), and domain-specific synthetic question-answer pairs.
@@ 5.1 Experimental Setup
Our evaluation spans four challenging benchmarks that test different aspects of ranking capability. BEIR Thakur et al. (2021) represents the gold standard for English retrieval evaluation, encompassing 13 heterogeneous tasks from question answering on Natural Questions to fact verification on FEVER, testing the model’s ability to generalize across domains without task-specific optimization. MIRACL Zhang et al. (2023b) pushes multilingual boundaries with 18 languages spanning diverse linguistic families, from Arabic and Chinese to Finnish and Thai, requiring deep cross-lingual understanding. MKQA Longpre et al. (2021) specifically challenges cross-lingual question answering capabilities, while CoIR Li et al. (2025) focuses on the specialized domain of code retrieval, where semantic understanding of programming constructs becomes crucial.
The first-stage dense retriever is jina-embeddings-v3, providing the foundation top-100 candidates that all rerankers process. Second-stage rerankers encompass our previous jina-reranker-v2, the multilingual bge-reranker-v2-m3, the mxbai-rerank variants at different scales, and Qwen3-Reranker-0.6B and Qwen3-Reranker-4B models.
@@ 5.2 Overall Performance Across Benchmarks
Table 2 demonstrates jina-reranker-v3’s exceptional performance density across diverse evaluation scenarios. On BEIR, our model achieves the highest score among all rerankers at 61.85, establishing new state-of-the-art performance for English retrieval. This represents a 4.79% improvement over our previous jina-reranker-v2 at 57.06, directly attributable to LBNL interaction mechanism where query and document embeddings are extracted from shared forward passes rather than separate encoding pipelines.
Parameter efficiency analysis reveals striking advantages compared to larger alternatives. Against the 1.5B parameter mxbai-rerank-large-v2, jina-reranker-v3 achieves superior BEIR performance with 61.85 versus 61.44 using $2.5\times$ fewer parameters, while providing specialized domain coverage unavailable in competing models reaching 70.64 on CoIR. This efficiency derives from architectural innovations: Qwen3’s optimized transformer backbone combined with our specialized 512-dimensional projector network that concentrates ranking signals without requiring massive parameter scaling.
Multilingual evaluation reveals strong cross-lingual capabilities despite the model’s compact architecture. The 66.83 score on MIRACL, while 2.49 points below the multilingual-specialized bge-reranker-v2-m3 at 69.32, demonstrates effective knowledge transfer from our progressive training methodology. The 67.92 MKQA performance closely approaches jina-reranker-m0’s 68.19, indicating that architectural sophistication can partially offset parameter differences in multilingual scenarios.
@@
jina-code-embeddings-0.5b - - - - 73.94
@@
jina-reranker-v3 0.6B 61.85 66.83 67.92 70.64
jina-reranker-v2 0.3B 57.06 63.65 67.90 58.35
jina-reranker-m0 2.4B 58.95 66.75 68.19 66.89
bge-reranker-v2-m3 0.6B 56.51 69.32 67.88 36.28
mxbai-rerank-base-v2 0.5B 58.40 55.32 64.24 65.71
mxbai-rerank-large-v2 1.5B 61.44 57.94 67.06 70.87
Qwen3-Reranker-0.6B 0.6B 56.28 57.70 65.34 65.18
Qwen3-Reranker-4B 4.0B 61.16 67.52 69.25 73.91
Table 2: Evaluation results for all rerankers. All scores are from our runs based on the top-100 retrieval results from the first row. For MKQA, we used Recall@10; for all other benchmarks, we used NDCG@10.
@@ 5.3 English Retrieval Performance on BEIR
Table 3 provides granular analysis across BEIR’s heterogeneous tasks, revealing specific architectural advantages. The model achieves consistent excellence across diverse reasoning tasks, with particularly strong performance on complex multi-hop reasoning reaching 78.58 on HotpotQA and fact verification achieving 94.01 on FEVER. These results highlight how LBNL interaction enables sophisticated query-document self-attention during encoding, capturing evidence relationships that separate encoding approaches miss.
Within the same scale category, jina-reranker-v3 reveals significant advantages. Against bge-reranker-v2-m3 with the same 0.6B parameters, jina-reranker-v3 delivers a substantial 5.34% improvement from 56.51 to 61.85, demonstrating architectural innovation over simple parameter scaling. The specialized 512-dimensional projector network effectively concentrates ranking signals while preserving contextual representations from the Qwen3 backbone. Remarkably, our model surpasses mxbai-rerank-large-v2’s 61.44 performance while using $2.5\times$ fewer parameters, establishing that sophisticated architecture can surpass brute-force scaling approaches.
Since jina-reranker-v3 processes all documents simultaneously in a listwise manner within shared context windows, we investigate the sensitivity to document ordering. We evaluate three variants: documents ordered by descending relevance scores (D), ascending scores (A), and random permutation (R). The results show modest variations across orderings, with random ordering (R) achieving the highest average of 62.24, followed by descending (D) at 61.85 and ascending (A) at 61.45. While the differences are not conclusive, this analysis reveals that the LBNL interaction maintains relatively stable performance across different input orderings, suggesting robust self-attention mechanisms that can effectively process documents regardless of their initial arrangement.
The model’s dominance extends particularly to question-answering scenarios, where Natural Questions achieves 74.28 and argumentative retrieval on ArguAna reaches 73.43, showcasing the benefit of contextual embeddings. These tasks require understanding complex query intent and matching it against nuanced document semantics, precisely the scenario where our dual embedding extraction approach provides maximum advantage over traditional cross-encoder scoring.
@@
jina-reranker-v3 (D) 0.6B 61.85 84.75 37.66 74.28 78.58 49.16 73.43 32.24 47.98 23.23 94.01 41.63 76.51 90.63
jina-reranker-v3 (A) 0.6B 61.45 85.90 39.14 72.34 77.48 50.99 69.36 29.73 48.30 23.90 93.46 41.72 76.75 89.73
jina-reranker-v3 (R) 0.6B 62.24 86.59 38.92 72.90 78.03 51.81 74.12 30.12 48.37 24.26 93.84 43.05 76.84 90.24
@@ 5.4 Multilingual Performance on MIRACL
Table 4: Multilingual retrieval performance on the MIRACL (measured by nDCG@10).
@@ 6 Conclusion
We present jina-reranker-v3, a 0.6B-parameter multilingual listwise reranker that introduces last but not late interaction for efficient document reranking. Our approach enables cross-document interactions during encoding by processing queries and multiple documents simultaneously within shared context windows, then extracting contextual embeddings from designated special token positions. By adapting long-context generative LLMs into a discriminative model, jina-reranker-v3 bridges the efficiency-effectiveness gap while maintaining significant computational advantages over generative rerankers. Future work includes studying the robustness of ranking against prompt injections and deduplication within the context window using submodularity optimization.